# Riccardo Giordana Pozzi

### v1

@@
We study universal features of defect correlation functions in supersymmetric defect CFTs, focusing on four-point functions of the displacement supermultiplet. By perturbing the leading-order correlators at strong coupling, we identify constraints on the operators exchanged at next-to-leading order. We then discuss the conditions under which correlation functions of identical superdisplacement components receive contributions only from multiparticle operators built from the same components. This leads to the equivalence between four-point functions of defect insertions across different theories, establishing universality. We confirm these features in examples including the $1/2$ BPS line in $\mathcal{N}=4$ SYM, $\mathcal{N}=2$ gauge theories and ABJM, up to the first subleading order in the strong-coupling expansion. We then analyze the $1/2$ BPS Wilson line in 3d $\mathcal{N}=2$ Chern–Simons matter theories, identifying the preserved defect algebra and the superdisplacement multiplet. Exploiting universality, we find the superdisplacement the four-point functions of the defect operators and we extract the conformal data up to the first subleading order. Finally, we comment on the $1/3$ BPS Wilson line in ABJM, providing initial results for the analysis at strong coupling.
@@ 1 Introduction and summary
The study of Conformal Field Theories (CFTs) in the presence of extended defects has developed over the years into a thriving field, providing a powerful framework for modeling systems with applications across many areas of physics. Defects, which may appear as dynamical extended operators, generically break part of the symmetry of the bulk CFT while preserving a residual conformal subgroup. This structure defines a defect Conformal Field Theory (dCFT) [1, 2], describing the localized degrees of freedom on the defect and their interactions with the bulk CFT. The discussion is further enriched when supersymmetry is present. Superconformal defects, which preserve a portion of the bulk supercharges, exhibit enhanced protection of their data, such as operator dimensions and OPE coefficients, and impose stronger constraints on their correlators.
@@
One of the most notable realizations of a line defect is the Wilson line in supersymmetric gauge theories [3, 4, 5, 6, 7], which provides a powerful laboratory for probing non-perturbative dynamics at strong coupling. In fact, when a holographic dual is available, correlation functions of local operator insertions along the line can be computed using Witten diagrams [8]. A well-known example is the $1/2$ BPS Wilson line in $\mathcal{N}=4$ Super Yang–Mills (SYM), where the Bremsstrahlung function of the theory is determined by the coefficient of the two-point functions of protected operator insertions on the Wilson line [9, 7]. Further insight into these operators has been gained through the computation of their four-point functions [8]. In the strong coupling regime, these correlation functions can be evaluated via the AdS/CFT correspondence [10]. Here, the dual string model corresponds to an effective field theory on $\mathrm{AdS}_{2}$, enabling the calculation of correlation functions using Witten diagram [8].
@@
However, a holographic realization of Wilson lines is not always available, and even in cases where it is known, computations via Witten diagrams quickly become demanding at higher orders. To address the limitations of standard perturbative holographic computations, which depend on detailed knowledge of the dual theory, the conformal bootstrap program [11, 12] has been extended to $1d$ dCFTs [13]. By constructing an ansatz and enforcing constraints such as crossing symmetry, internal consistency and mild growth conditions on anomalous dimensions at large scaling dimensions [14, 15, 16], one gains access to correlation functions at strong coupling. Moreover, this approach can be algorithmically applied to study higher-order subleading corrections.
@@
The $1/2$ BPS Wilson line in ${\mathcal{N}}=4$ SYM was the first case studied through an analytical bootstrap approach, carried out up to next-to-leading order (NLO) in the strong-coupling expansion [14]. This laid the groundwork for a more detailed understanding of Wilson lines in this theory [17, 18, 19] and paved the way for numerous applications to other defects, including the $1/2$ BPS Wilson line in ${\mathcal{N}}=2$ gauge theories [20]. Then, the same framework has been extended to lower-dimensional settings, such as the $1/2$ BPS Wilson line in ABJM theory [21] and in ${\mathcal{N}}=4$ Chern–Simons-matter (CSm) theories [22], which naturally serve as lower-dimensional analogues of the higher-dimensional examples.
@@
Evidence that this general structure persists at subleading orders comes from the study of the dual string theory realizations of these setups. In particular, when analyzing the bulk string description of Wilson lines via Witten diagrams up to NLO, one finds that the structures arising from the tree-level expansion of the Nambu-Goto action are equivalent. This is, for example, the case for the $1/2$ BPS Wilson line in SYM and ABJM, [8, 21]. A concrete example is provided by the four-point function of the orthogonal fluctuations, which are dual to displacement insertions. In these cases, the tree-level interactions never mix with the elementary excitations associated with fluctuations along the compact manifold. This shows that different theories produce equivalent structures, i.e. equivalent four-point functions, see section 5.
@@ Results
Most of our analysis relies on the observation that, when perturbing around a given leading-order theory, the spectrum of operators that can be exchanged at the first subleading order can be constrained by studying mixed correlation functions. Consider, schematically, the $1d$ mixed four-point function with two generic insertions, $\mathbb{D}$ and $\mathbb{O}$. At NLO one finds
$\epsilon\,f_{\mathbb{D}\mathbb{D}\mathbb{O}\mathbb{O}}^{(1)}(z)=\sum_{\Delta}(\lambda_{\Delta}^{(0)}\tilde{\lambda}_{\Delta}^{(1)}+\lambda_{\Delta}^{(1)}\tilde{\lambda}_{\Delta}^{(0)})g_{\Delta}(z)+\lambda_{\Delta}^{(1)}\tilde{\lambda}_{\Delta}^{(1)}g_{\Delta}(z)+\lambda_{\Delta}^{(0)}\tilde{\lambda}_{\Delta}^{(0)}\gamma^{(1)}_{\Delta}(g_{\Delta}(z)\log(-z)+\mathrm{g}_{\Delta}(z))\,$,
arising from perturbing the leading order conformal data $(\lambda_{\Delta}^{(0)},\tilde{\lambda}_{\Delta}^{(0)},\Delta)$ with the NLO corrections $(\lambda_{\Delta}^{(1)},\tilde{\lambda}_{\Delta}^{(1)},\gamma_{\Delta}^{(1)})$. Here, $g_{\Delta}(z)$ denotes the $1d$ conformal block, $\mathrm{g}_{\Delta}(z)$ a specific derivative of the block with respect to the conformal dimension, arising from the expansion, and $\epsilon f_{\mathbb{D}\mathbb{D}\mathbb{O}\mathbb{O}}^{(1)}(z)$ the first-order correction to the correlator. The OPE coefficients are concisely denoted by
$\lambda_{\Delta}^{(i)}=\lambda_{\mathbb{D}\mathbb{D}\mathbb{T}_{\Delta}}^{(i)}\qquad\text{and}\qquad\tilde{\lambda}_{\Delta}^{(i)}=\lambda_{\mathbb{O}\mathbb{O}\mathbb{T}_{\Delta}}^{(i)}\,$,
with $\mathbb{T}_{\Delta}$ the exchanged primary of dimension $\Delta$.
The expression can be interpreted as a constraint on the minimal order in the $\epsilon$ expansion of the OPE data. For instance, suppose one of the two leading-order coefficients vanishes, e.g. $\lambda_{\Delta^{\star}}^{(0)}=0$ while $\tilde{\lambda}_{\Delta^{\star}}^{(0)}\neq 0$. In this case, consistency requires $\lambda_{\Delta^{\star}}^{(1)}\sim\mathcal{O}(\epsilon)$, obtained by matching the expansion orders on both sides. This constraint becomes particularly significant when studying the four-point function with all $\mathbb{D}$ insertions, where at NLO one finds
@@
$\epsilon\,f_{\mathbb{D}\mathbb{D}\mathbb{D}\mathbb{D}}^{(1)}(z)=\sum_{\Delta}(\lambda_{\Delta}^{(0)}\lambda_{\Delta}^{(1)}+\lambda_{\Delta}^{(1)}\lambda_{\Delta}^{(0)})g_{\Delta}(z)+\lambda_{\Delta}^{(1)}\lambda_{\Delta}^{(1)}g_{\Delta}(z)+\lambda_{\Delta}^{(0)}\lambda_{\Delta}^{(0)}\gamma^{(1)}_{\Delta}(g_{\Delta}(z)\log(-z)+\mathrm{g}_{\Delta}(z))\,.$
When we look at the values for the exchanged contributions associated to $\lambda_{\Delta^{\star}}^{(0)}=0$ and $\tilde{\lambda}_{\Delta^{\star}}^{(0)}\neq 0$ only the quadratic combination $\lambda_{\Delta^{\star}}^{(1)}\lambda_{\Delta^{\star}}^{(1)}$ survives. However, being this of order ${\mathcal{O}}(\epsilon^{2})$ we can conclude that the corresponding operator is not exchanged up to NLO, see subsection 2.2 for a more detailed discussions.
This line of reasoning allows us to assess the presence of universal features in a given system. Under additional hypotheses, e.g. on the spectrum, one can show that four-point functions built from repeated insertions of a given multiplet component do not exchange operators from different families up to NLO. In this sense, the correlator factorizes: its conformal block decomposition involves only operators within the universal sector, without mixing from other exchanges. With additional assumptions on the OPE of the supermultiplet, such four-point functions can thus be identified with universal sectors.
@@
Beyond providing the physical insight of uncovering such features, the main advantage of identifying universal sectors is that it allows for the determination of a four-point function up to NLO directly, bypassing the standard bootstrap strategy of postulating and constraining an ansatz. Once universal sectors have been identified, subsequent analyses can build on established results rather than rederiving them from first principles. This way, it is possible to obtain four-point functions even in systems that cannot be fully constrained, either due to limited supersymmetry or an insufficient set of constraints. Moreover, relying on previous results where the expansion parameter $\epsilon$ was fixed in terms of the two-point normalization of the supermultiplet, either explicitly by matching the holographic theories [14, 21] or just by CFT based arguments [24, 25], one can incorporate this relation into the discussion of universal sectors, thereby determining the same expression for other cases as well. Consequently, universal four-point functions share equivalent expressions up to NLO.
@@
We verify these arguments by showing that universal features are indeed present in a variety of known defect theories. When the defects have the same codimension, the identification for the four-point function of the displacement is straightforward. All the others follows by requiring supersymmetry. For instance, the displacement four-point functions of the $1/2$ BPS line in ABJM [21] and in $\mathcal{N}=4$ CSm theories [22] fall within the same universality class. Up to NLO one find the expression
@@
$+$ $\Bigl(6-\frac{16}{z}+6z^{4}-16z^{5}\Bigr)\log(1-z)+2z^{4}(8z-3)\log(-z)\Bigr)\Bigr]\,$
where $z$ is a cross-ratio, $C_{\mathcal{T}}$ and $\mathbb{n}_{\mathcal{T}}$ are theory-specific constants, related to the Bremsstrahlung function and to the preserved supersymmetry, respectively. By $\mathcal{T}$} we mean the $1/2$ BPS line ether in ABJM or in ${\mathcal{N}}=4$ CSm theories.
@@
$\langle\mathbb{D}^{i}(t_{1}){\mathbb{D}}^{j}(t_{2})\mathbb{D}^{k}(t_{3}){\mathbb{D}}^{l}(t_{4})\rangle_{{\mathcal{W}}_{\text{SYM}}}=\frac{{\mathcal{C}}_{\text{SYM}}^{2}}{t_{12}^{4}t_{34}^{4}}G^{ijkl}(z)$
@@
$G^{ijkl}(z)=$ $\Bigl[\delta_{ij}\delta_{kl}\,G_{S}^{(0)}(z)+(\delta_{ij}\delta_{kl}-\delta_{il}\delta_{jk})G_{A}^{(0)}(z)+(\delta_{ik}\delta_{jl}+\delta_{il}\delta_{jk}-2/3\,\delta_{ij}\delta_{kl})G_{T}^{(0)}(z)+$
$+$ $\frac{1}{4\pi^{2}C_{\text{SYM}}}\Bigl(\delta_{ij}\delta_{kl}\,G_{S}^{(1)}(z)+(\delta_{ij}\delta_{kl}-\delta_{il}\delta_{jk})G_{A}^{(1)}(z)+(\delta_{ik}\delta_{jl}+\delta_{il}\delta_{jk}-2/3\,\delta_{ij}\delta_{kl})G_{T}^{(1)}(z)\Bigr)\Bigr]\,.$
By properly dealing with the tensorial structure one obtains an expression for the displacement insertions in the ABJM Wilson line given by
@@
which, when explicitly evaluated, matches with the previous one and with the results of [21].
Although the above arguments were illustrated for the displacement operator, they extend to any four-point function of the supermultiplet components. An especially interesting case arises when considering the four-point function of the superprimary of the $1/2$ BPS Wilson line in ${\mathcal{N}}=2$ gauge theories in $4d$, as discussed in [20]. In this setup, universality does not extend to the entire displacement supermultiplet, but one can still identify subsectors in which it applies. This is the case, for instance, for the superprimary four-point function. By showing that this correlator does not receive contributions from other sectors up to NLO, we are able to identify it explicitly. Furthermore, we find consistency with the results of [20] by fixing the free parameters left undetermined in the bootstrap approach to $c_{1}=c_{2}=-4$, thereby fully determining the superprimary four-point function to be
@@
$+(-4+2z+z^{3}-4z^{4})\frac{\log(1-z)}{4z}+\frac{z^{2}(2-4z+9z^{2}-7z^{3}+2z^{4})}{2(z-1)^{3}}\log(-z)\Bigr)\Bigr]\,$,
and fixing the anomalous dimensions [20] to
@@
In a similar way, we identify a universal sector for the $1/2$ BPS line in $AdS_{3}\times S^{3}\times T^{4}$ [26, 27] with purely Ramond–Ramond flux, finding consistency with previous analysis and fixing the overall two-point normalization of the displacement supermultiplet in terms of the bulk coupling.
We then provide brief general arguments supporting universality from the holographic perspective.
In line with previous analyses of $1/2$ BPS Wilson lines at strong coupling in $3d$ SCFTs, such as ABJM [21] and ${\mathcal{N}}=4$ CSm theories [22], we investigate the case of ${\mathcal{N}}=2$ CSm theories. We examine the preserved defect algebra, identify the superdisplacement multiplet ($\mathcal{D}$}), and employ superspace methods to express the four-point functions in terms of cross-ratios. We study the OPE structure for the superdisplacement in the two different channels: the chiral–chiral ${\mathcal{D}}\times{\mathcal{D}}$ channel, which contains one short multiplets and infinitely many long operators with unprotected, higher dimensions, and the chiral–antichiral ${\mathcal{D}}\times\bar{\mathcal{D}}$ channel, where only long multiplets appear [21, 28]. We thus derive the corresponding superblocks by diagonalizing the superconformal Casimir.
By identifying the universal sectors, we determine the four-point functions of the superdisplacement up to leading and next-to-leading order at strong coupling. For the superprimary we find
@@
We conclude this work with preliminary results for the $1/3$ BPS Wilson line in ABJM, discussed in detail in [29], at strong coupling. In particular, we argue that universal features can also be identified in this case, allowing for the determination of the four-point functions of the displacement and tilt (tlit) supermultiplets, respectively. The analysis of the displacement will be equivalent to that of the Wilson line in ${\mathcal{N}}=4$ CSm theory. Therefore, we focus on the tilt multiplet which four-point function is specified by that of the superprimary
@@
with $n$ label of the GFF multiparticle states $\mathbbm{F}\bar{\mathbbm{F}}]_{n}^{L=2}$.
@@
While universality is not expected to hold beyond NLO, it is still plausible that higher-order corrections to four-point functions contain universal contributions mixed with theory-specific ones. In such cases, one can decompose a given correction into a universal part and a non-universal remainder. Identifying and isolating these universal sectors can provide valuable input for bootstrapping higher-order four-point functions.
@@
Finally, while the present discussion has focused primarily on Wilson lines and line defects, the underlying philosophy of constraining the order of possible exchanges is more general and can be generalized and applied to higher-dimensional defects, as for example surface defects [30], or perhaps even to strongly coupled CFTs beyond defect theories.

### v2

@@
We study universal features of defect correlation functions in supersymmetric defect CFTs, focusing on four-point functions of the displacement supermultiplet. By perturbing the leading-order correlators at strong coupling, we derive constraints that identify the operators exchanged at next-to-leading order. From this, we determine the conditions under which four-point functions of defect insertions across different theories are equivalent, thereby establishing universality. We confirm these features in examples including the $1/2$ BPS line in $\mathcal{N}=4$ SYM, $\mathcal{N}=2$ gauge theories and ABJM, up to the first subleading order in the strong-coupling expansion. We further analyze the $1/2$ BPS line in 3d $\mathcal{N}=2$ Chern-Simons-matter theories, identifying the preserved defect algebra and the superdisplacement multiplet. Exploiting universality, we find the superdisplacement four-point functions of defect operators and extract the conformal data at first subleading order. Finally, we comment on the $1/3$ BPS Wilson line in ABJM, presenting initial results from the strong-coupling analysis.
@@ 1 Introduction and summary
The study of conformal field theories (CFTs) in the presence of extended defects has developed over the years into a thriving field, providing a powerful framework for modeling systems with applications across many areas of physics. Defects, which may appear as dynamical extended operators, generically break part of the symmetry of the bulk CFT while preserving a residual conformal subgroup. This structure defines a defect conformal field theory (dCFT) [9, 21], describing the localized degrees of freedom on the defect and their interactions with the bulk CFT. The discussion is further enriched when supersymmetry is present. Superconformal defects, which preserve a portion of the bulk supercharges, exhibit enhanced protection of their data, such as operator dimensions and operator product expansion (OPE) coefficients, and impose stronger constraints on their correlators.
@@
One of the most notable realizations of a line defect is the Wilson line in supersymmetric gauge theories [48, 20, 22, 54, 18], which provides a powerful laboratory for probing nonperturbative dynamics at strong coupling. In fact, when a holographic dual is available, correlation functions of local operator insertions along the line can be computed using Witten diagrams [41]. A well-known example is the $1/2$ BPS Wilson line in $\mathcal{N}=4$ super-Yang-Mills (SYM), where the bremsstrahlung function of the theory is determined by the coefficient of the two-point functions of protected operator insertions on the Wilson line [15, 18]. Further insight into these operators has been gained through the computation of their four-point functions [41]. In the strong coupling regime, these correlation functions can be evaluated via the AdS/CFT correspondence [47]. Here, the dual string model corresponds to an effective field theory on $\mathrm{AdS}_{2}$, enabling the calculation of correlation functions using Witten diagrams [41].
@@
However, a holographic realization of Wilson lines is not always available, and even in cases where it is known, computations via Witten diagrams quickly become demanding at higher orders. To address the limitations of standard perturbative holographic computations, which depend on detailed knowledge of the dual theory, the conformal bootstrap program [53, 31] has been extended to $1d$ dCFTs [32]. By constructing an ansatz and enforcing constraints such as crossing symmetry, internal consistency, and mild growth conditions on anomalous dimensions at large scaling dimensions [46, 43, 37], one gains access to correlation functions at strong coupling. Moreover, this approach can be algorithmically applied to study higher-order subleading corrections.
@@
The $1/2$ BPS Wilson line in ${\mathcal{N}}=4$ SYM was the first case studied through an analytical bootstrap approach, carried out up to next-to-leading order (NLO) in the strong-coupling expansion [46]. This laid the groundwork for a more detailed understanding of Wilson lines in this theory [33, 34, 35] and paved the way for numerous applications to other defects, including the $1/2$ BPS Wilson line in ${\mathcal{N}}=2$ gauge theories [40]. Then, the same framework was extended to lower-dimensional settings, such as the $1/2$ BPS Wilson line in ABJM theory [7] and in ${\mathcal{N}}=4$ Chern-Simons-matter (CSm) theories [52], which naturally serve as lower-dimensional analogs of the higher-dimensional examples.
@@
Evidence that this general structure persists at subleading orders comes from the study of the dual string theory realizations of these setups. In particular, when analyzing the bulk string description of Wilson lines via Witten diagrams up to NLO, one finds that the structures arising from the tree-level expansion of the Nambu-Goto action are equivalent. This is, for example, the case for the $1/2$ BPS Wilson line in SYM and ABJM, [41, 7]. A concrete example is provided by the four-point function of the orthogonal fluctuations, which are dual to displacement insertions. In these cases, the tree-level interactions never mix with the elementary excitations associated with fluctuations along the compact manifold. This shows that different theories produce equivalent structures, i.e. equivalent four-point functions, see Sec. 5.
@@ Results
Most of our analysis relies on the observation that, when perturbing around a given leading-order theory, the spectrum of operators that can be exchanged at the first subleading order can be constrained by studying mixed correlation functions. We briefly introduce here the line of reasoning that is discussed completely in Sec. 2.2. In this regard, we consider a CFT1 toy model with two protected operators, that we call $\mathbb{D}(t)$ and $\mathbb{O}(t)$, and towers of higher-particle operators built by these fundamental words, e.g. $\mathbb{D}\mathbb{D}]^{L=2}_{n}$. As we will see these will play the role of defect operators. We are interested in studying the four-point function of a given operator such as
$\langle\mathbb{D}(t_{1})\mathbb{D}(t_{2})\mathbb{D}(t_{3})\mathbb{D}(t_{4})\rangle=\frac{C^{2}}{t_{12}^{2\Delta_{\mathbb{D}}}t_{34}^{2\Delta_{\mathbb{D}}}}f_{\mathbb{D}^{4}}(z)\,$,
where $C$ is an overall normalization constant and $z=\frac{t_{12}t_{34}}{t_{14}t_{32}}$ is the cross-ratio. We also assume, for later convenience, that the function $f(z)$ is expressed as an expansion about a given leading-order theory by
$f(z)=f_{\mathbb{D}^{4}}^{(0)}(z)+\epsilon\,f_{\mathbb{D}^{4}}^{(1)}(z)+{\mathcal{O}}(\epsilon^{2})\,$,
where $f^{(0)}(z)$ represent the leading-order theory and $f^{(1)}(z)$ represent a small deviation from that. In the following, we take $f^{(0)}(z)$ to be that of a GFF, which thus allows us to consider
$f_{\mathbb{D}^{4}}^{(0)}(z)=\sum_{\Delta}(\lambda^{(0)}_{\Delta})^{2}\,g_{\Delta}(z)\,$,
where $g_{\Delta}(z)$ are the $1d$ conformal blocks and $\lambda^{(0)}_{\Delta}=\lambda^{(0)}_{\mathbb{D}\mathbb{D}\mathbb{T}_{\Delta}}$ denote the OPE coefficients associated with the exchanged operators, here generically called $\mathbb{T}_{\Delta}$. The spectrum of operators that can appear at this order can be easily identified by performing Wick contractions, which show that only $\mathbb{D}\mathbb{D}]^{L=2}_{n}$ operators are exchanged at leading-order. One can then move to the more interesting case in which interactions are included. In this case, the OPE coefficients receive next-to-leading-order corrections parametrized by $\lambda^{(1)}_{\Delta}$ and the conformal dimensions acquire anomalous contributions $\gamma_{\Delta}$. The conformal block expansion of $f^{(1)}(z)$ can therefore be obtained by perturbing the GFF expansion, leading to
@@
$\epsilon\,f_{\mathbb{D}^{4}}^{(1)}(z)=\sum_{\Delta}\Bigl(2\lambda_{\Delta}^{(0)}\lambda_{\Delta}^{(1)}+(\lambda_{\Delta}^{(1)})^{2}\Bigr)g_{\Delta}(z)+(\lambda_{\Delta}^{(0)})^{2}\,\gamma^{(1)}_{\Delta}(g_{\Delta}(z)\log(-z)+\mathrm{g}_{\Delta}(z))\,$,
with $\mathrm{g}_{\Delta}(z)$ representing a particular derivative contribution of the conformal block $g_{\Delta}(z)$, (2.14).
However, at this order the identification of the exchanged operator is not straightforward. In this respect, the issue can be addressed by analyzing mixed four-point functions and reinterpreting the resulting expressions as constraints on the order in $\epsilon$ of the OPE coefficients. We introduce this idea by studying the $1d$ mixed four-point function of $\mathbb{D}$ and $\mathbb{O}$. At NLO one finds
$\epsilon\,f_{\mathbb{D}^{2}\mathbb{O}^{2}}^{(1)}(z)=\sum_{\Delta}(\lambda_{\Delta}^{(0)}\tilde{\lambda}_{\Delta}^{(1)}+\lambda_{\Delta}^{(1)}\tilde{\lambda}_{\Delta}^{(0)})g_{\Delta}(z)+\lambda_{\Delta}^{(1)}\tilde{\lambda}_{\Delta}^{(1)}g_{\Delta}(z)+\lambda_{\Delta}^{(0)}\tilde{\lambda}_{\Delta}^{(0)}\gamma^{(1)}_{\Delta}(g_{\Delta}(z)\log(-z)+\mathrm{g}_{\Delta}(z))\,$,
arising from perturbing the leading order conformal data. In order to distinguish the OPE coefficients, we further introduced $\tilde{\lambda}^{(i)}_{\Delta}=\lambda_{\mathbb{O}\mathbb{O}\mathbb{T}_{\Delta}}^{(i)}$. As anticipated, this expression can be read as a constraint on the OPE data, allowing us to identify the minimum order in $\epsilon$ at which these contributions can appear consistently with the expansion.
For instance, consider a single contribution associated with a particular $\mathbb{T}_{\Delta^{\star}}$ such that the two leading OPE data satisfy $\lambda_{\Delta^{\star}}^{(0)}=0$ while $\tilde{\lambda}_{\Delta^{\star}}^{(0)}\neq 0$. This is the case, for example, for the multiparticle operators $\mathbb{O}\mathbb{O}]^{L=2}_{n}$. In this situation, consistency requires $\lambda_{\Delta^{\star}}^{(1)}\sim\mathcal{O}(\epsilon)$, as follows from matching the orders in $\epsilon$ on both sides of the relation. We can thus, in a sense, bootstrap the order of the individual OPE coefficients.
Although this conclusion may appear innocuous, this information becomes particularly valuable when inserted into the expansion of $\epsilon\,f_{\mathbb{D}^{4}}^{(1)}(z)$ discussed above. This is because, given the hypothesis $\lambda_{\Delta^{\star}}^{(0)}=0$, the only contribution that can appear associated to the exchange of $\mathbb{T}_{\Delta^{\star}}$ is the quadratic term $(\lambda_{\Delta^{\star}}^{(1)})^{2}$. Since we have identified $\lambda_{\Delta^{\star}}^{(1)}\sim\mathcal{O}(\epsilon)$, such contributions are of order $\mathcal{O}(\epsilon^{2})$ and therefore do not appear up to NLO. See Sec. 2.2 for a more detailed discussions.
This approach provides a powerful tool for identifying universal properties within a given system. At the same time, it is not restricted to the specific setup considered here and can be easily extended and applied to other contexts. We therefore expect it to remain applicable in more general settings and to provide useful insights elsewhere as well, as discussed later. In the present case, we apply these methods to supersymmetric defect CFTs at strong coupling. Under additional assumptions, one can show that four-point functions built from repeated insertions of a given component of a supermultiplet do not exchange operators from different families up to NLO.
In this sense, the correlator factorizes: in a four-point function with identical external operators—such as four tilt or four displacement operators—the exchanged spectrum consists only of multiparticle operators constructed from the external one.
Consider, for example, two defect operators with the same protected scaling dimension in two different theories. If their four-point functions receive no contributions from multiparticle operators or from fundamental operators associated with other fields in the respective supermultiplet, the resulting four-point structures would be equivalent, provided these have similar three-point interactions. In this sense, the corresponding correlation functions are universal. See Sec. 3 for the full discussion.
@@
Beyond providing the physical insight of uncovering such features, the main advantage of identifying universal sectors is that it allows for the determination of a four-point function up to NLO directly, bypassing the standard bootstrap strategy of postulating and constraining an ansatz. Once universal sectors have been identified, subsequent analyses can build on established results rather than rederiving them from first principles. This way, it is possible to obtain four-point functions even in systems that cannot be fully constrained, either due to limited supersymmetry or an insufficient set of constraints. For recent developments in this regards see [42, 23].
Moreover, identifying the parameter $\epsilon$ is in general a complicated task, as it cannot be determined by bootstrap arguments alone. Its identification can therefore be achieved only by explicitly matching to holographic theories, when available [46, 7], or alternatively through purely CFT-based arguments [25, 38]. In the present case, we exploit the identification of the universal sector to determine the form of the parameter $\epsilon$ by relying on previously studied examples in which it has already been fixed. Indeed, once a correlation function is identified in a given case, all correlators within the same universal class can be straightforwardly determined. In this way, the problem of explicitly computing these quantities is reduced to demonstrating that two different theories belong to the same universal class.
@@
We verify these arguments by showing that universal features are indeed present in a variety of known defect theories. When the defects have the same codimension, the identification for the four-point function of the displacement is straightforward. All the others follows by requiring consistency with supersymmetry. For instance, the displacement four-point functions of the $1/2$ BPS line in ABJM [7] and in $\mathcal{N}=4$ CSm theories [52] fall within the same universality class, as shown in Sec. 4.1. Up to NLO one finds the universal expression
@@
$+$ $\Bigl(6-\frac{16}{z}+6z^{4}-16z^{5}\Bigr)\log(1-z)+2z^{4}(8z-3)\log(-z)\Bigr)\Bigr]\,$,
where now $\mathbb{D}(t)$ is the actual displacement operator while $C_{\mathcal{T}}$ and $\mathbb{n}_{\mathcal{T}}$ are theory-specific constants, related to the particular bremsstrahlung function and to the preserved supersymmetry, respectively. By $\mathcal{T}$} we mean the $1/2$ BPS line ether in ABJM or in ${\mathcal{N}}=4$ CSm theories.
@@
$\langle\mathbb{D}^{i}(t_{1}){\mathbb{D}}^{j}(t_{2})\mathbb{D}^{k}(t_{3}){\mathbb{D}}^{l}(t_{4})\rangle_{{\mathcal{W}}_{\text{SYM}}}=\frac{{\mathcal{C}}_{\text{SYM}}^{2}}{t_{12}^{4}t_{34}^{4}}G^{ijkl}(z)\,$,
@@
$G^{ijkl}(z)=$ $\Bigl[\delta^{ij}\delta^{kl}\,G_{S}^{(0)}(z)+(\delta^{ij}\delta^{kl}-\delta^{il}\delta^{jk})G_{A}^{(0)}(z)+(\delta^{ik}\delta^{jl}+\delta^{il}\delta^{jk}-2/3\,\delta^{ij}\delta^{kl})G_{T}^{(0)}(z)+$
$+$ $\frac{1}{4\pi^{2}C_{\text{SYM}}}\Bigl(\delta^{ij}\delta^{kl}\,G_{S}^{(1)}(z)+(\delta^{ij}\delta^{kl}-\delta^{il}\delta^{jk})G_{A}^{(1)}(z)+(\delta^{ik}\delta^{jl}+\delta^{il}\delta^{jk}-2/3\,\delta^{ij}\delta^{kl})G_{T}^{(1)}(z)\Bigr)\Bigr]\,.$
By properly dealing with the tensorial structure, i.e. taking the embedding, one obtains an expression for the displacement insertions in the ABJM Wilson line given by
@@
which, when explicitly evaluated, agrees with the previous result as well as with the findings of [7], as discussed in more details in Sec. 4.2.
Although the above arguments were illustrated for the displacement operator, they extend to any four-point function of the supermultiplet components. An especially interesting case arises when considering the four-point function of the superprimary of the $1/2$ BPS Wilson line in ${\mathcal{N}}=2$ gauge theories in $4d$, discussed in [40]. In this setup, universality does not extend to the entire displacement supermultiplet, but one can still identify subsectors in which it applies. In Sec. 4.3, we show that this is the case for the superprimary four-point function. By showing that this correlator does not receive contributions from other sectors up to NLO, we are able to identify it explicitly. Furthermore, we find consistency with the results of [40] by fixing the free parameters left undetermined in the bootstrap approach to
$c_{1}=c_{2}=-4\,$,
thereby fully determining the superprimary four-point function to be
@@
$+(-4+2z+z^{3}-4z^{4})\frac{\log(1-z)}{4z}+\frac{z^{2}(2-4z+9z^{2}-7z^{3}+2z^{4})}{2(z-1)^{3}}\log(-z)\Bigr)\Bigr]\,.$
Through these explicit values, we finally fix the anomalous dimensions [40] to
@@
In Sec. 4.4, we use a similar approach to identify a universal sector for the $1/2$ BPS line in $AdS_{3}\times S^{3}\times T^{4}$ [10, 14] with purely Ramond-Ramond flux, finding consistency with previous analysis and fixing the overall two-point normalization of the displacement supermultiplet in terms of the bulk coupling.
We then provide brief general arguments supporting universality from the holographic perspective, in Sec. 5.
In line with previous analyses of $1/2$ BPS Wilson lines at strong coupling in $3d$ SCFTs, such as ABJM [7] and ${\mathcal{N}}=4$ CSm theories [52], we investigate the case of ${\mathcal{N}}=2$ CSm theories, in Sec. 6. We examine the preserved defect algebra, identify the superdisplacement multiplet ($\mathcal{D}$}), and employ superspace methods to express the four-point functions in terms of cross-ratios. We study the OPE structure for the superdisplacement in the two different channels: the chiral-chiral ${\mathcal{D}}\times{\mathcal{D}}$ channel, which contains one short multiplet and infinitely many long operators with unprotected, higher dimensions, and the chiral-antichiral ${\mathcal{D}}\times\bar{\mathcal{D}}$ channel, where only long multiplets appear [7, 8]. We thus derive the corresponding superblocks by diagonalizing the superconformal Casimir.
By identifying the universal sectors, we determine the four-point functions of the superdisplacement up to leading- and next-to-leading order at strong coupling. For the superprimary we find
@@
Sec. 7 concludes this work with preliminary results for the $1/3$ BPS Wilson line in ABJM at strong coupling, discussed in detail in [26]. In particular, we argue that universal features can also be identified in this case, allowing for the determination of the four-point functions of the displacement and tilt (tlit) supermultiplets, respectively. The analysis of the displacement will be equivalent to that of the Wilson line in ${\mathcal{N}}=4$ CSm theory. Therefore, we focus on the tilt multiplet which four-point function is specified by that of the superprimary
@@
with $n$ the label of the GFF multiparticle states $\mathbbm{F}\bar{\mathbbm{F}}]_{n}^{L=2}$.
@@
While universality is not expected to hold beyond NLO, it is still plausible that higher-order corrections to four-point functions contain universal contributions mixed with theory-specific ones. In such cases, one can decompose a given correction into a universal part and a nonuniversal remainder. Identifying and isolating these universal sectors can provide valuable input for bootstrapping higher-order four-point functions.
@@
Finally, while the present discussion has focused primarily on Wilson lines and line defects, the underlying philosophy of constraining the order of possible exchanges is more general and can be generalized and applied to higher-dimensional defects, as, for example, surface defects [19], or perhaps even to strongly coupled CFTs beyond defect theories.